# MUSE spectroscopy of the high abundance discrepancy planetary nebula NGC 6153

### v2

@@ Abstract
Results. Our analysis confirmed the presence of a low-temperature plasma component in NGC 6153. We find that electron temperatures derived from recombination line/continuum diagnostics are significantly lower than those derived from collisionally excited lines diagnostics. Ionic chemical abundance maps were constructed, considering the weight of the cold plasma phase in the H i emission. Adopting this approach we found ionic abundances that could be up to 0.2 dex lower for those derived from CELs and up to 1.1 dex higher for those derived from RLs than in the case of an homogeneous H i emission. The Abundance Contrast Factor (ACF) between both plasma components was defined, with values, on average, 0.9 dex higher than the ADF. Different methods for calculating ionisation correction factors (ICFs), including state-of-the-art literature ICFs and machine learning techniques, yielded consistent results.
Conclusions. Our findings emphasize that accurate chemical abundance determinations in high-ADF PNe must account for multiple plasma phases. Future research should focus on expanding this methodology to a broader sample of PNe, with spectra deep enough to gather physical condition information of both plasma components, which will enhance our understanding of their chemical compositions and the underlying physical processes in these complex objects.
@@
One of the unsolved questions in nebular astrophysics is whether the correct determination of the elemental abundances of elements heavier than helium in photoionised nebulae is provided by bright collisionally excited lines (CELs) or by faint recombination lines (RLs, see e. g. García-Rojas & Esteban, 2007). In almost all objects studied so far, this abundance discrepancy is characterised for a given ion by the ratio of abundances obtained from RLs and CELs, which is called the abundance discrepancy factor (ADF). The ADF values are greater than 1, with a median value of about 2–3 for H II regions and the majority of planetary nebulae (PNe).111For an updated compilation of literature values of the ADF(O2+) see https://nebulousresearch.org/adfs/. Several scenarios have been proposed to explain this behaviour over the past decades. The most popular ones include the presence of temperature inhomogeneities in the gas (Peimbert, 1967; Torres-Peimbert et al., 1980), chemical inhomogeneities in the gas (Torres-Peimbert et al., 1990; Liu et al., 2000; Tsamis & Péquignot, 2005), or deviations from a Maxwellian distribution of the energy of thermal electrons (kappa distribution; Nicholls et al., 2012). Obtaining direct evidence to support any of these scenarios has proven to be a challenging task. However, observational evidence of the action of temperature fluctuations in photoionised nebulae has been found very recently, although restricted to H ii regions (Méndez-Delgado et al., 2023).
@@
Most of the scenarios proposed to explain the AD fail when the ADF is too high. This is the case with the increasing number of PNe exhibiting ADF values much higher than those observed in H ii regions, reaching values of 10 or more. However, several studies have proposed that the presence of at least two distinct plasma components in these PNe is the driver of the high ADFs found. Although this scenario was proposed more than thirty years ago to explain the general behaviour of AD in ionised nebulae (Torres-Peimbert et al., 1990), it has only gained relevance in the last decade due to increasing observational evidence of the presence of two plasma components (e.g., Corradi et al., 2015; García-Rojas et al., 2016, 2022; Richer et al., 2022). In this scenario, one of the plasma components is very metal-rich (i.e. hydrogen-poor) and exhibits very different physical conditions compared to the main gas component that emits CELs. The origin of this metal-rich gas component is still a matter of debate, and several scenarios have been proposed, most of them linked to the fact that the majority of the high-ADF PNe appear to be associated with a close binary central star that has undergone a common envelope event (see e.g. Corradi et al., 2015; Jones et al., 2016; Wesson et al., 2018).
@@
In this context, in order to understand the physical processes that give rise to the phenomenon of high-ADF PNe, we need to refine the abundance determinations for each of the components. The advent of the Multi Unit Spectroscopic Explorer (MUSE) integral-field spectrograph (Bacon et al., 2010) on the Very Large Telescope (VLT) has revolutionised the field of PN research, paving the way for high-spatial resolution studies which allow detailed mapping of the physical and chemical properties on these objects. However, the first studies of the PNe NGC 7009, NGC 3242 and IC 418 (Walsh et al., 2016, 2018; Monreal-Ibero & Walsh, 2020; Walsh & Monreal-Ibero, 2020) were done using the nominal mode, which covers the wavelength range 470930 nm that prevents the detection of the O ii RLs around 465 nm and hence, not allowing to address the AD problem. In order to circumvent this issue, García-Rojas et al. (2022, hereafter GR22) obtained deep observations of three high-ADF PNe using the MUSE extended mode, which covers the wavelength range nm, and mapping, for the first time, the spatial distribution of the O ii RLs in this type of PNe. GR22 emphasised the importance of taking into account the contribution of recombination to the [O ii] and [N ii] auroral lines, which are crucial for a proper determination of the physical conditions of the warm gas, and the relative importance of hydrogen emission in both components, which is essential to properly compute the metal content of the cold gas component, as was pointed out in the photoionisation models presented by Gómez-Llanos & Morisset (2020). Regarding this last issue, GR22 presented a methodology to compute the relative contribution of H to the H-poor gas component; however, given the lack of diagnostics to estimate the physical conditions of the H-poor gas in the three PNe studied, only rough estimates could be made in that study.
@@
NGC 6153 is a bright, southern PN, with a relatively high ADF10 (Liu et al., 2000). The large amount of deep spectrophotometric data obtained for this object has made it the most studied PN in terms of the AD problem. Several observational techniques, from moderate-spectral-resolution longslit spectroscopy scanning the entire nebular surface (Liu et al., 2000) to high-spectral-resolution échelle spectroscopy (McNabb et al., 2016; Richer et al., 2022), have been used to analyse the physical conditions and chemical abundances in this object.
@@ 2 Observations and data reduction
NGC 6153 was observed with MUSE on the VLT, in seeing-limited mode, on the night of 6 to 7 July 2016. The log of the observations is provided in Table 1. The sky conditions were reasonable with a seeing of around 1 arcsec and under thin clouds. The instrument was used in its Wide Field Mode with the natural seeing (WFM–NOAO) configuration. This provides a nearly contiguous 1 arcmin2 field of view with 0.2-arcsec spatial sampling. We used the extended mode of MUSE (WFM–NOAO-E), which covers the wavelength range nm with an effective spectral resolution that increases from at the bluest wavelengths to at the reddest wavelengths. The data are available at the ESO Science Archive under Prog. ID 097.D–0241(A) (PI: R. L. M. Corradi). Observations were obtained with dithering and a rotation of 90° between each exposure to remove artifacts during data processing. Given that the target is extended, we obtained two sky frames on each sequence (see Table 1) to perform an adequate sky subtraction. A few short exposures were also taken to analyse the strongest emission lines. The data reduction was done with esorex, using the dedicated ESO pipeline (Weilbacher et al., 2014, 2020). For each individual frame the pipeline performs bias subtraction, flat-fielding and slice-tracing, wavelength calibration, geometric corrections, illumination correction using twilight sky flats, sky subtraction (telluric absorption/emission lines and continuum) making use of the sky frames obtained between science exposures. In the final steps of the reduction process, differential atmospheric correction and flux calibration were performed.
@@
![Refer to caption](x1.png)
As in GR22, we checked for possible saturation of the brightest lines with a preliminary inspection of the data cubes with qfitsview (Ott, 2012), finding that [O iii]5007 and H lines were slightly saturated in the brightest zones. We therefore avoid using [O iii]5007 measurements and to minimize the effects of saturation in H, we followed the same procedure as in GR22 to compute the extinction coefficient.
@@ 3 Emission line measurements
We constructed flux maps for 60 emission lines (30 RLs and 30 CELs). The complete list of emission lines is shown in Table C. All the fluxes and associated errors were measured as described in Sect. 3 of GR22. In addition, we constructed continuum maps at 8100 Å and 8400 Å, using the automated line-fitting algorithm ALFA (Wesson, 2016). To only consider high signal-to-noise ratio observations, we define the following mask: F(H)0.005 F(H)max, that we apply to all emission line maps and their subsequent analysis.
In Fig. 16 we present the unreddened fluxes of the most relevant emission lines, sorted by atomic mass and ionisation potential (IP) of the ion. The RLs of metals present a similar spatial distribution regardless of the element and IP of the emitting ion (see C ii6462, N ii5679, N iii4641222Line N iii4640.64 is blended with O ii4641.81, and possibly with O ii4638.86 and N iii4641.85., O i7773+, and O ii4649333Corresponds to O ii4649.13 + 4650.85, and possibly affected by a blend with C iii4650.25. We will check the effect of this blend in chemical abundance maps in Sect. 4.5.), unlike the emission maps of CELs where there is a clear ionisation structure when comparing the emission of ions with different IP of the same or different element (e.g., [N i]5198, [N ii]6548, [Ar iii]7136, [Ar iv]4740, [Ar v]7005). The emission line [O ii]7330+, which shows a morphology similar to the RLs of metals in the inner parts of the nebula, has a non-negligible recombination contribution that is estimated and subtracted from the total emission in Sec.4.3.
We present a spatially-resolved map of a neutron-capture emission line ([Kr iv] at 5867.74 Å) (see Fig. 16). To our knowledge, this emission has been previously spatially mapped only in the PN IC 2165 Otsuka (2022). The [Kr iv]5867.74 line emission map clearly resembles the spatial distribution of emission lines with a similar ionisation potential range, such as [O iii]4959.
@@ 4.1 Presence of a metal-rich, cold plasma phase
As already mentioned in Sect. 1, in PNe demonstrating high ADFs (greater than 10), the presence of a high-metallicity central cold phase of gas is now well established, particularly owing to the efforts made by the community in obtaining spatially resolved observations of these types of objects (Wesson et al., 2003; Corradi et al., 2015; García-Rojas et al., 2016, 2022; Richer et al., 2022). In the subsequent sections of this paper, we will compute the ionic and total abundances accounting for the existence of two phases of gas: one being the “classical” warm shell emitting the CELs, and the other, more centrally emitted, metal-rich (or H-poor, with the difference lying in the He abundance; see below) and cold, as metal IR lines efficiently cool the gas. This second component serves as the primary source of the metal RLs.
@@
In the following sections, we will operate under the assumption that the emission lines resulting from collisional excitation originate exclusively from the warm region, while those resulting from electron recombination to a metal ion originate exclusively from the cold region. Some lines may arise from both processes (e.g., auroral lines); this will be discussed in Section 4.3.
@@ 4.2 Extinction correction
The observed flux map of H is presented in Fig. 2, in logarithmic scale. From this bright emission line, we can distinguish the main nebular shell with two bright regions located northwest and southeast of the central star and a diffuse envelope.
![Refer to caption](x2.png)
![Refer to caption](x3.png)
![Refer to caption](x4.png)
@@
Ueta & Otsuka (2021) state that for spatially resolved observations, the spatial variations in c(H) cannot be recovered using a uniform (, ). An iteration is performed between the spaxel-by-spaxel values of and determined later in the process, and c(H), for different diagnostic line ratios. We calculate the distribution of c(H) using five different combinations of spatially resolved and . In Fig. 3 we present these five c(H) distributions as well as the c(H) based on uniform = 3,400 cm-3 and = 8,000 K, and several estimates from the literature444Kingsburgh & Barlow (1994), Liu et al. (2000), Pottasch et al. (2003), Tsamis et al. (2003), McNabb et al. (2016), and R22. The c(H) distributions are very consistent with each other and with the values from the literature, except for Kingsburgh & Barlow (1994) and R22, which show much lower extinctions, especially R22, whose c(H) is completely outside the distributions computed here.
Given the similarities of the results presented in Fig. 3 and to avoid adding noise to all emission maps, we choose to use c(H) obtained from uniform values for and instead of spatially resolved maps.
The resulting map of c(H) is shown in Fig. 4. We notice spatial variations in the extinction map on the order of approximately 6 per cent around a median value of 1.2, with a pattern similar to the spatial distribution of F(H) (see Fig. 2). This indicates that part of the extinction is due to dust located inside the nebula. A similar behaviour is found for NGC 7009 (Walsh et al., 2016), where c(H) roughly follows F(H). These authors compute the extinction map based on MUSE observations of the same Balmer and Paschen emission line ratios, correcting H and H from the He ii Pickering lines and following an iterative process to recalculate the theoretical line ratios based on the latter estimations of ([S iii]) and ([Cl iii]). The variations in c(H) found for NGC 7009 are much higher (more than 30 per cent) than what we find here for NGC 6153.
The presence of the cold component described in Sec. 4.1 can affect the determination of the extinction correction because it affects the theoretical value to which the observed value is compared to obtain the correction. Anticipating the results presented in the following sections, we take both components into account using Eq. 6 from in Morisset et al. (2023) and compute the theoretical value for H/H as being between 2.90 and 3.03, the minimum being the case of a single region at = 3,400 cm-3 and = 8,000 K, and the maximum corresponding to spaxels where the contribution of the cold region is maximum. These high values in the theoretical H/H would reduce the determination of c(H) from e.g. 1.2 to 1.14 at maximum. This would, for example, reduce the [S iii] ([N ii]) electron temperatures from 8,300 K to 8,247 K (8,278 K respectively). Therefore, to avoid adding noise to the data, we do not apply any correction for the presence of the cold region.

### v3

@@ Abstract
Results. Our analysis confirmed the presence of a low-temperature plasma component in NGC 6153. We find that electron temperatures derived from recombination line and continuum diagnostics are significantly lower than those derived from collisionally excited line diagnostics. Ionic chemical abundance maps were constructed, considering the weight of the cold plasma phase in the H i emission. Adopting this approach we found ionic abundances that could be up to 0.2 dex lower for those derived from CELs and up to 1.1 dex higher for those derived from RLs than in the case of a homogeneous H i emission. The abundance contrast factor (ACF) between both plasma components was defined, with values, on average, 0.9 dex higher than the ADF. Different methods for calculating ionisation correction factors (ICFs), including state-of-the-art literature ICFs and machine learning techniques, yielded consistent results.
Conclusions. Our findings emphasise that accurate chemical abundance determinations in high-ADF PNe must account for multiple plasma phases. Future research should focus on expanding this methodology to a broader sample of PNe, with spectra deep enough to gather physical condition information of both plasma components, which will enhance our understanding of their chemical compositions and the underlying physical processes in these complex objects.
@@
One of the unsolved questions in nebular astrophysics is whether the correct determination of the elemental abundances of elements heavier than helium in photoionised nebulae is provided by bright collisionally excited lines (CELs) or by faint recombination lines (RLs, see e.g. García-Rojas & Esteban, 2007). In almost all objects studied so far, this abundance discrepancy is characterised for a given ion by the ratio of abundances obtained from RLs and CELs, which is called the abundance discrepancy factor (ADF). The ADF values are greater than 1, with a median value of about 2–3 for H II regions and the majority of planetary nebulae (PNe).111For an updated compilation of literature values of the ADF(O2+), see https://nebulousresearch.org/adfs/. Several scenarios have been proposed to explain this behaviour over the past decades. The most popular ones include the presence of temperature inhomogeneities in the gas (Peimbert, 1967; Torres-Peimbert et al., 1980), chemical inhomogeneities in the gas (Torres-Peimbert et al., 1990; Liu et al., 2000; Tsamis & Péquignot, 2005), or deviations from a Maxwellian distribution of the energy of thermal electrons (kappa distribution; Nicholls et al., 2012). Obtaining direct evidence to support any of these scenarios has proven to be a challenging task. However, observational evidence of the action of temperature fluctuations in photoionised nebulae has been found very recently, although it is restricted to H ii regions (Méndez-Delgado et al., 2023).
@@
Most of the scenarios proposed to explain the AD fail when the ADF is too high. This is the case with the increasing number of PNe exhibiting ADF values much higher than those observed in H ii regions, reaching values of 10 or more. However, several studies have proposed that the presence of at least two distinct plasma components in these PNe is the driver of the high ADFs found. Although this scenario was proposed more than thirty years ago to explain the general behaviour of AD in ionised nebulae (Torres-Peimbert et al., 1990), it has only gained relevance in the last decade due to increasing observational evidence of the presence of two plasma components (e.g. Corradi et al., 2015; García-Rojas et al., 2016, 2022; Richer et al., 2022). In this scenario, one of the plasma components is very metal-rich (i.e. hydrogen-poor) and exhibits very different physical conditions compared to the main gas component that emits CELs. The origin of this metal-rich gas component is still a matter of debate, and several scenarios have been proposed, most of them linked to the fact that the majority of the high-ADF PNe appear to be associated with a close binary central star that has undergone a common envelope event (see e.g. Corradi et al., 2015; Jones et al., 2016; Wesson et al., 2018).
@@
In this context, in order to understand the physical processes that give rise to the phenomenon of high-ADF PNe, we need to refine the abundance determinations for each of the components. The advent of the Multi Unit Spectroscopic Explorer (MUSE) integral-field spectrograph (Bacon et al., 2010) on the Very Large Telescope (VLT) has revolutionised the field of PN research, paving the way for high-spatial resolution studies that allow for detailed mapping of the physical and chemical properties on these objects. However, the first studies of the PNe NGC 7009, NGC 3132, and IC 418 (Walsh et al., 2016, 2018; Monreal-Ibero & Walsh, 2020, 2022) were done using the nominal mode, which covers the wavelength range 470930 nm that prevents the detection of the O ii RLs around 465 nm and hence, not allowing the AD problem to be addressed. In order to circumvent this issue, García-Rojas et al. (2022, hereafter GR22) obtained deep observations of three high-ADF PNe using the MUSE extended mode, which covers the wavelength range nm, and mapping, for the first time, the spatial distribution of the O ii RLs in this type of PN. GR22 emphasised the importance of taking into account the contribution of recombination to the [O ii] and [N ii] auroral lines, which are crucial for a proper determination of the physical conditions of the warm gas, and the relative importance of hydrogen emission in both components, which is essential to properly compute the metal content of the cold gas component, as was pointed out in the photoionisation models presented by Gómez-Llanos & Morisset (2020). Regarding this last issue, GR22 presented a methodology to compute the relative contribution of H to the H-poor gas component; however, given the lack of diagnostics to estimate the physical conditions of the H-poor gas in the three PNe studied, only rough estimates could be made in that study.
@@
NGC 6153 is a bright, southern PN, with a relatively high ADF10 (Liu et al., 2000). The large amount of deep spectrophotometric data obtained for this object has made it the most studied PN in terms of the AD problem. Several observational techniques, from moderate-spectral-resolution long-slit spectroscopy scanning the entire nebular surface (Liu et al., 2000) to high-spectral-resolution échelle spectroscopy (McNabb et al., 2016; Richer et al., 2022), have been used to analyse the physical conditions and chemical abundances in this object.
@@ 2 Observations and data reduction
NGC 6153 was observed with MUSE on the VLT, in seeing-limited mode, on the night of 6 to 7 July 2016. The log of the observations is provided in Table 1. The sky conditions were reasonable with a seeing of around 1 arcsec and under thin clouds. The instrument was used in its Wide Field Mode with the natural seeing (WFM–NOAO) configuration. This provides a nearly contiguous 1 arcmin2 field of view with 0.2-arcsec spatial sampling. We used the extended mode of MUSE (WFM–NOAO-E), which covers the wavelength range nm with an effective spectral resolution that increases from at the bluest wavelengths to at the reddest wavelengths. The data are available at the ESO Science Archive under Prog. ID 097.D–0241(A) (PI: R. L. M. Corradi). Observations were obtained with dithering and a rotation of 90° between each exposure to remove artefacts during data processing. Given that the target is extended, we obtained two sky frames on each sequence (see Table 1) to perform an adequate sky subtraction. A few short exposures were also taken to analyse the strongest emission lines. The data reduction was done with esorex, using the dedicated ESO pipeline (Weilbacher et al., 2014, 2020). For each individual frame the pipeline performs bias subtraction, flat-fielding and slice-tracing, wavelength calibration, geometric corrections, illumination correction using twilight sky flats, sky subtraction (telluric absorption/emission lines and continuum) making use of the sky frames obtained between science exposures. In the final steps of the reduction process, differential atmospheric correction and flux calibration were performed.
@@
As in GR22, we checked for possible saturation of the brightest lines with a preliminary inspection of the data cubes with qfitsview (Ott, 2012), finding that [O iii]5007 and H lines were slightly saturated in the brightest zones. We therefore avoid using [O iii]5007 measurements and to minimise the effects of saturation in H, we followed the same procedure as in GR22 to compute the extinction coefficient.
@@ 3 Emission line measurements
We constructed flux maps for 60 emission lines (30 RLs and 30 CELs). The complete list of emission lines is shown in Table 9. All the fluxes and associated errors were measured as described in Sect. 3 of GR22. In addition, we constructed continuum maps at 8100 Å and 8400 Å, using the automated line-fitting algorithm ALFA (Wesson, 2016). To only consider high signal-to-noise ratio observations, we define the following mask: F(H)0.005 F(H)max, that we apply to all emission line maps and their subsequent analysis.
In Fig. 16 we present the unreddened fluxes of the most relevant emission lines, sorted by atomic mass and ionisation potential (IP) of the ion. The RLs of metals present a similar spatial distribution regardless of the element and IP of the emitting ion (see C ii6462, N ii5679, N iii4641222Line N iii4640.64 is blended with O ii4641.81, and possibly with O ii4638.86 and N iii4641.85., O i7773+, and O ii4649333Corresponds to O ii4649.13 + 4650.85, and possibly affected by a blend with C iii4650.25. We will check the effect of this blend in chemical abundance maps in Sect. 4.5.), unlike the emission maps of CELs where there is a clear ionisation structure when comparing the emission of ions with different IP of the same or different element (e.g. [N i]5198, [N ii]6548, [Ar iii]7136, [Ar iv]4740, [Ar v]7005). The emission line [O ii]7330+, which shows a morphology similar to the RLs of metals in the inner parts of the nebula, has a non-negligible recombination contribution that is estimated and subtracted from the total emission in Sec.4.3.
We present a spatially resolved map of a neutron-capture emission line ([Kr iv] at 5867.74 Å) (see Fig. 16). To our knowledge, this emission has been previously spatially mapped only in the PN IC 2165 Otsuka (2022). The [Kr iv]5867.74 line emission map clearly resembles the spatial distribution of emission lines with a similar ionisation potential range, such as [O iii]4959.
@@ 4.1 Presence of a metal-rich, cold plasma phase
As already mentioned in Sect. 1, in PNe demonstrating high ADFs (greater than 10), the presence of a high-metallicity central cold phase of gas is now well established, particularly owing to the efforts made by the community in obtaining spatially resolved observations of these types of objects (Wesson et al., 2003; Corradi et al., 2015; García-Rojas et al., 2016, 2022; Richer et al., 2022). In the subsequent sections of this paper, we explain how we computed the ionic and total abundances accounting for the existence of two phases of gas: one being the “classical” warm shell emitting the CELs, and the other, more centrally emitted, metal-rich (or H-poor, with the difference lying in the He abundance; see below) and cold, as metal IR lines efficiently cool the gas. This second component serves as the primary source of the metal RLs.
@@
In the following sections, we will operate under the assumption that the emission lines resulting from collisional excitation originate exclusively from the warm region, while those resulting from electron recombination to a metal ion originate exclusively from the cold region. Some lines may arise from both processes (e.g. auroral lines); this will be discussed in Section 4.3.
@@ 4.2 Extinction correction
The observed flux map of H is presented in Fig. 2, in logarithmic scale. From this bright emission line, we can distinguish the main nebular shell with two bright regions located north-west and south-east of the central star and a diffuse envelope.
@@
Ueta & Otsuka (2021) state that for spatially resolved observations, the spatial variations in c(H) cannot be recovered using a uniform (, ). An iteration is performed between the spaxel-by-spaxel values of and determined later in the process, and c(H), for different diagnostic line ratios. We calculate the distribution of c(H) using five different combinations of spatially resolved and . In Fig. 3 we present these five c(H) distributions as well as the c(H) based on uniform = 3,400 cm-3 and = 8,000 K, and several estimates from the literature444Kingsburgh & Barlow (1994), Liu et al. (2000), Pottasch et al. (2003), Tsamis et al. (2003), McNabb et al. (2016), and R22. The c(H) distributions are very consistent with each other and with the values from the literature, except for Kingsburgh & Barlow (1994) and R22, which show much lower extinctions, especially R22, whose c(H) is completely outside the distributions computed here. Given the similarities of the results presented in Fig. 3 and to avoid adding noise to all emission maps, we choose to use c(H) obtained from uniform values for and instead of spatially resolved maps.
The resulting map of c(H) is shown in Fig. 4. We notice spatial variations in the extinction map of the order of approximately 6 per cent around a median value of 1.2, with a pattern similar to the spatial distribution of F(H) (see Fig. 2). This indicates that part of the extinction is due to dust located inside the nebula. A similar behaviour is found for NGC 7009 (Walsh et al., 2016), where c(H) roughly follows F(H). These authors compute the extinction map based on MUSE observations of the same Balmer and Paschen emission line ratios, correcting H and H from the He ii Pickering lines and following an iterative process to recalculate the theoretical line ratios based on the latter estimations of ([S iii]) and ([Cl iii]). The variations in c(H) found for NGC 7009 are much higher (more than 30 per cent) than what we find here for NGC 6153.
The presence of the cold component described in Sec. 4.1 can affect the determination of the extinction correction because it affects the theoretical value to which the observed value is compared to obtain the correction. Anticipating the results presented in the following sections, we take both components into account using Eq. 6 from in Morisset et al. (2023) and compute the theoretical value for H/H as being between 2.90 and 3.03, the minimum being the case of a single region at = 3,400 cm-3 and = 8,000 K, and the maximum corresponding to spaxels where the contribution of the cold region is maximum. These high values in the theoretical H/H would reduce the determination of c(H) from, for example, 1.2 to 1.14 at maximum. This would, for example, reduce the [S iii] ([N ii]) electron temperatures from 8,300 K to 8,247 K (8,278 K respectively). Therefore, to avoid adding noise to the data, we do not apply any correction for the presence of the cold region.